# ©Plug-in Authorization for Human Content Copyright Protection in Text-to-Image Model

### v1

@@ A.2 Technique Related: Concept Removal, and Negative Sampling
Our extraction operation is closely related with the concept removal for generative models. [26, 25] remove target concepts by matching the generation distribution of contexts with target concepts and that of contexts without target concepts. [46] forget target concepts by minimizing the cross attention of target concepts with that of target images. Heng and Soh [47] leverage the reverse process of continual learning to promote the controllable forgetting of target contents in deep generative models.
@@
We note that negative sampling [22] can also prevent generating certain concepts. Specifically, end users can set conditional context and negative context to guide the diffusion process to generate images conforming the conditional context while being far away from the negative context. Only negative sampling cannot stop copyright infringing generation because the contexts are set freely and adversarially by end users.
@@
In contrast, for a specific copyrighted concept, our extraction operation takes an “Reverse LoRA” approach to disentangle the base model into two parts: the non-infringing base model and the plug-in LoRA component for the copyrighted concept. Specifically, we use negative sampling to generate non-infringing images, which serve as training data for copyright plug-in. From the aspect of parameter efficient fine-tuning, our paper is related to literature [48, 49, 50, 10, 51].
@@
Our combination operation is related with the widely studied “knowledge distillation” [52, 53, 54, 55, 56], but entails large difference from previous work. We combine multiple copyright plug-ins that are LoRA components for different targets, and we take data free approach due to practical constraint.
Appendix B Intermediate results of extraction
@@
Appendix C Experiment on ordinary objects generation
We evaluated the influence of extraction on the generation of ordinary objects. We utilize 5000 textual captions selected from the validation set in MS-COCO [57] as prompts, generating 5000 images using SD1.5 and the non-infringing model that extracts R2D2 and Picasso, respectively. Several randomly selected images are displayed in Figure 6. For illustrative purposes, we also generated results for concept-ablation and ESD on MS-COCO, respectively. Images within the same column exhibit substantial similarity, indicating that extraction does not exert an impact on the generation of ordinary items.
@@ Appendix C Experiment on ordinary objects generation
Figure 6: Ordinary objects generation after extraction. Row 1 displays images generated by Stable Diffusion V-1.5. Rows 2 and 3 illustrate images generated after the removal of the IP character R2D2, while Rows 4 and 5 showcase images generated after the elimination of Picasso’s style. Rows 3 and 5 serve as the baseline, representing concept-ablation and ESD, respectively. Notably, after the extraction of R2D2 and Picasso, the non-infringing model retains the capability to generate commonplace objects sourced from the MS-COCO dataset [57].
@@
As depicted in Table3, we calculate some quantitative metrics like FID and KID. It is noteworthy that Concept-Ablation [25] only releases the checkpoint of ablating “R2D2” and ESD [26] only releases the checkpoint of erasing “Picasso”. Thus, we compare them separately by using their respective checkpoints. Evaluation metrics consistently maintain low values, further affirming that extraction does not compromise the generation of ordinary items.

### v2

@@ 4.2 Technique Related: Concept Removal, and Negative Sampling
Our extraction operation is closely related with the concept removal for generative models. Rohit Gandikota and Joanna Materzyńska and Jaden Fiotto-Kaufman and David Bau (2023); Kumari et al. (2023) remove target concepts by matching the generation distribution of contexts with target concepts and that of contexts without target concepts. Zhang et al. (2023a) forget target concepts by minimizing the cross attention of target concepts with that of target images. Heng & Soh (2023) leverage the reverse process of continual learning to promote the controllable forgetting of target contents in deep generative models.
@@
We note that negative sampling (Ho & Salimans, 2022) can also prevent generating certain concepts. Specifically, end users can set conditional context and negative context to guide the diffusion process to generate images conforming the conditional context while being far away from the negative context. Only negative sampling cannot stop copyright infringing generation because the contexts are set freely and adversarially by end users.
@@
In contrast, for a specific copyrighted concept, our extraction operation takes an “Reverse LoRA” approach to disentangle the base model into two part: non-infringing base model and the plug-in LoRA component for copyrighted concept. Specifically, we use negative sampling to generate non-infringing images, which serves as training data for copyright plug-in. From the aspect of parameter efficient fine-tuning, our paper is related with literature (Alaluf et al., 2022; Ruiz et al., 2023; Gal et al., 2022; Hu et al., 2022; Huang et al., 2023).
@@
Our combination operation is related with the widely studied “knowledge distillation” (Liang et al., 2023; Lopes et al., 2017; Sun et al., 2019; Hinton et al., 2015; Fang et al., 2019), but entails large difference from previous work. We combine multiple copyright plug-ins that are LoRA components for different targets, and we take data free approach due to practical constraint.
5 Discussion, Open Questions and Limitations
The growing concerns regarding generative AI models stem from their capacity to produce copyright-infringing content. This issue becomes more pronounced as state-of-the-art models continue to improve the quality of generated images, often without adequately acknowledging the contributions of human content creators. In response, we propose the “Copyright Plug-in Authorization” framework to address these societal worries, drawing inspiration from the purpose of copyright law. Our approach demonstrates that copyrighted data can be incorporated into LoRA plug-ins, enabling straightforward tracking of usage and equitable distribution of rewards.
A key challenge for this framework is the efficient management of a large number of plug-ins, which is essential to ensure user-friendly access to specific generations. Moreover, updating the base model poses another challenge, as retraining the entire suite of plug-ins can be costly, raising the issue of ensuring backward compatibility. One limitation of our current research is the potential degradation in the performance of the non-infringing model due to a large number of extraction operations, a factor that has yet to be thoroughly investigated.
Appendix A Appendix
A.1 Intermediate results of extraction
@@
A.2 Experiment on ordinary objects generation
We evaluated the influence of extraction on the generation of ordinary objects. We utilize 5000 textual captions selected from the validation set in MS-COCO (Lin et al., 2015) as prompts, generating 5000 images using SD1.5 and the non-infringing model that extracts R2D2 and Picasso, respectively. Several randomly selected images are displayed in Figure 8. For illustrative purposes, we also generated results for concept-ablation and ESD on MS-COCO, respectively. Images within the same column exhibit substantial similarity, indicating that extraction does not exert an impact on the generation of ordinary items.
@@ A.2 Experiment on ordinary objects generation
Figure 8: Ordinary objects generation after extraction. Row 1 displays images generated by Stable Diffusion V-1.5. Rows 2 and 3 illustrate images generated after the removal of the IP character R2D2, while Rows 4 and 5 showcase images generated after the elimination of Picasso’s style. Rows 3 and 5 serve as the baseline, representing concept-ablation and ESD, respectively. Notably, after the extraction of R2D2 and Picasso, the non-infringing model retains the capability to generate commonplace objects sourced from the MS-COCO dataset (Lin et al., 2015).
@@
As depicted in Table3, we calculate some quantitative metrics like FID and KID. It is noteworthy that Concept-Ablation (Kumari et al., 2023) only releases the checkpoint of ablating “R2D2” and ESD (Rohit Gandikota and Joanna Materzyńska and Jaden Fiotto-Kaufman and David Bau, 2023) only releases the checkpoint of erasing “Picasso”. Thus, we compare them separately by using their respective checkpoints. Evaluation metrics consistently maintain low values, further affirming that extraction does not compromise the generation of ordinary items.
A.3 More Quantitative Results
We compare the results of three methods following the removal of Van Gogh’s influence. Like LPIPS, DINO-v2 Oquab et al. (2023) is a self-supervised model used for extracting visual features from images, allowing us to compute the similarity between generated images generated by the non-infringing model and the base model. The higher the value of DINO-v2, the closer the image is. Also, we calculate the CLIP-t distance to measure the semantic similarity between images generated by the non-infringing model and their corresponding textual prompts. The higher the CLIP-t value, the closer the sematic similarity is.
To ensure fairness in our evaluation, we implemented various settings to control the extraction effect and align it as closely as possible with the surrounding style. The quantitative results, as detailed in Table 4, demonstrate that our method achieves superior removal of the target style while preserving the integrity of the surrounding style.
Table 4: Quantitative comparison of different methods. The symbol $\uparrow$ indicates that the higher value is better on the metric, whereas $\downarrow$ symbol signifies the lower value is more preferable.The colors of symbols $\uparrow$ and $\downarrow$ should be read according to the colors of the most left column (Style).
[…]
Target Style ESD 138 0.385 46.06 26.9
EXTRACTION (OURS) 187 0.387 38.8 23.9
Surrounding Style ESD 27 0.212 71.42 32.15
EXTRACTION (OURS) 32 0.157 79.02 31.9
Target Style CONCEPT-ABLATION 42 0.255 52.1 28.1
EXTRACTION (OURS) 58 0.293 48.6 28.2
Surrounding Style CONCEPT-ABLATION 12 0.123 81.2 29.6
EXTRACTION (OURS) 7.4 0.115 86.2 32.8
A.4 Experiment on seen and unseen contents generation
In extraction, we sample 10 common contents (training set) leveraging ChatGPT to fine-tune the base model. These contents have been previously processed by the non-infringing model. Additionally, we generate 10 supplementary contents for evaluation. Figure 9 shows the images generated with these 20 contents. The images on the left represent the seen contents (training set), while those on the right are the unseen contents (evaluation set). Within each image block, we extract the corresponding artistic style. The top row is generated by the base model, wheras the bottom row is generated by the non-infringing model. All images are generated with the prompt "[content] by [artist]". The specifics of the artistic and contents are detailed in Table 5.
Figure 9: Seen and Unseen contents generation after extraction. On the left are 10 contents already seen during the extraction, while on the right aren’t seen. Within each image block, the top row is generated by the base model, and the non-infringing model generates the bottom one. Zoom in for better visualization.
Table 5: Details of artistic style and contents
Artistic Style Seen Contents (training set) Unseen Contents (evaluation set)
Claude Monet vase of flowers barn in a rural setting
Vincent van Gogh bowl of fruit bustling street market
Pablo Picasso still life with candles boat on a calm river
Johannes Vermeer landscape with rolling hills group of animals in a field
Leonardo da Vinci cityscape with buildings crowded café scene
Michelangelo forest with sunlight filtering through trees horse grazing in a pasture
Frida Kahlo portrait of a person vintage clock on a mantelpiece
Rembrandt van Rijn quiet beach at sunset window with a view of the countryside
Salvador Dalí mountain range with snow room with antique furniture
Henri Matisse tranquil lake with reflections close-up of a tree’s bark and leaves
To evaluate the performance of the base model and the non-infringing model, we computed various image distance metrics, including KID, LPIPS, and DINO-v2, between images generated by these two models. Additionally, we assessed the CLIP-t distance to measure the semantic alignment between textual prompts and images produced by the non-infringing model.
The quantitative results, averaged on the seen and unseen contents separately, are summarized in Table 6. The lack of significant differences across all metrics for the seen versus unseen content suggests that the model did not exhibit signs of over-fitting.
Table 6: Quantitative results on seen and unseen contents. The gap in metrics between seen content and unseen content is not significant. The symbol $\uparrow$ indicates that the higher value is better on the metric, whereas $\downarrow$ symbol signifies the lower value is more preferable.
[…]
Target Style Seen Contents 137 0.315 46.8 28.5
Unseen Contents 168 0.303 43.3 28.0
Surrounding Style Seen Contents 17.92 0.139 81.0 32.3
Unseen Contents 16.66 0.135 81.6 32.1
A.5 Extracting specific prompt v.s. extracting style
Figure 10 compares the outputs of the base model and the non-infringing model under two different extraction scenarios, extracting specific prmpt "sunflowers by van gogh" and entire style "van gogh". For each scenario presented, the left column exhibits images generated by the base model, whereas the right column features images produced by the non-infringing model. The results indicate that the extraction method is effective not only in isolating and modifying individual content-style associations but also in comprehensively altering an artist’s entire style.
Table 7 shows that both extracting specific prompts and styles have little impact on the model’s generative ability.
Figure 10: The results of the non-infringing model after extracting specific prompt and style. In the left block, only the specific prompt "sunflowers by van gogh" is extracted, while in the right block the entire style "van gogh" is extracted. For each block, the left column displays images generated by the base model, whereas the right column shows images generated by the corresponding non-infringing model.
Table 7: Quantitative results on COCO caption set.
[…]
Extract "van gogh" 22 3.4
Extract "sunflower by van gogh" 20.8 1.4
A.6 Style-IP ©plug-in combination
Figure 11 shows the combination of style and IP ©plug-ins. The initial image is generated by the non-infringing model after extracting R2D2 and van Gogh, where the IP and style ate hard to recognize. Subsequent images, the second and third, are produced by the model after separately adding the R2D2 and van Gogh ©plug-in, thus distinctly featuring the respective IP and style. The final image is created after adding the combined ©plug-in, successfully achieving IP recreation and style replication.
Figure 11: IP Recreation and style replication in a single image. We can integrate ©plug-in into the non-infringing model to recreate R2D2 or replicate van Gogh’s style in a single image. Also, we can integrate the combined ©plug-in to achieve them both in an image.
A.7 The limitation of potential degradation in the performance of the non-infringing model
We observe a decline in the performance of the non-infringing model as the number of extracted styles increases. This degradation can be effectively mitigated by increasing the rate of re-context iterations and de-concept iterations, i.e., we can sample more contexts in the re-contexts in the re-context sub-process to maintain the generative ability of the non-infringing model.
During the extract process, we set the number of de-concept iterations to be 10 and set the number of re-context iterations to $10\times r$. The default value for the rate is 1. For extraction, we leverage ChatGPT to generate $10\times r$ common contents. For each iteration, we select one of these contents to generate 8 images. A higher $r$ allows the re-context phase to learn more contextual information, thereby maintaining the model’s ability to generate high-quality images even as the number of extracted styles grows.
Figure 12 illustrates this phenomenon. The first row displays the image generated by the base model. Subsequent rows, from top to bottom, are images generated by non-infringing models that extract 1 to 10 styles, respectively. As the number of extracted styles increases, there is a noticeable decline in image quality. However, by increasing the rate, this degradation is significantly alleviated, as evidenced by improved image fidelity. The styles and contexts are detailed in Table 5. When the rate is greater than 1, we use the context in the evaluation set for the re-context phase.
Figure 12: The degradation and alleviation in the performance of the non-infringing model. We randomly sample 8 images generated by different models. The first line is the image generated by the base model. The rest, from top to bottom, are images generated by non-infringing models that extract 1, 2, 3…, 10 styles, respectively. As the extracted concept number increases, the quality of images continuously declines. With the increase of the rate of re-context iteration number and de-concept iteration number, this decline has been effectively alleviated. Zoom in for better visualization.